A sample of 0.5moles of chlorine is in a 150ml container STP(standard temp and volume (0C and 101.3kPa). If 8.5 more moles of chlorine is added to the container, what would happen to the volume? Calculate the new volume using avogardos law. Show work and formula
A sample of chlorine is in a 150mL container at 22kPa with a temp of 20C. If the temp of the container is increased to 105C and the pressure increases to 250kPa. What would happen to the volume? Calculate the bew volume using combined gas law. Show work and formula.
1
Expert's answer
2019-05-29T05:16:35-0400

a) Using Avogadro's law "\frac{V_1}{n_1}=\frac{V_2}{n_2}" we can calculate volume after addition of chlorine:

"V_2=V_1 \cdot \frac{n_2}{n_1} = 0.15 L \cdot \frac{0.5 mol + 8.5 mol}{0.5 mol}=2.7L"

b) Using combined gas law "\frac{p_1V_1}{T_1}=\frac{p_2V_2}{T_2}" we can calculate volume after increasing of temperature and pressure.

"p_1=22kPa=22000Pa; V_1= 150mL = 0.15L; T_1=20^{\circ}C=293K"

"p_2=250kPa=250000Pa; T_2=105^{\circ}C=378K"

"V_2=V_1 \cdot \frac{p_1}{p_2} \cdot \frac {T_2}{T_1}=0.15L \cdot \frac{22000Pa}{250000Pa} \cdot \frac{378K}{293K}=0.017L=17mL"